OpenStudy (anonymous):

Can someone show me how to show that the sum of the divisors of 2^(k-1) is (2^(k+1) -1) I'm pretty sure you need to use the Proposition for Divisors from Prime Factorization. I need this for a proof I am doing. Thanks! Can someone show me how to show that the sum of the divisors of 2^(k-1) is (2^(k+1) -1) I'm pretty sure you need to use the Proposition for Divisors from Prime Factorization. I need this for a proof I am doing. Thanks! @Mathematics

OpenStudy (turingtest):

There aren't very many here who can do these kinds of problems, so look for Zarkon, JamesJ, or Harkirat. I've seen them do this kind of thing befor.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm pretty sure the sum of the divisors for 2^(k-1) looks like this \[\sum_{i=1}^{k - 1} 2i\]

OpenStudy (anonymous):

how can I ask them for help?

OpenStudy (turingtest):

Look for them on the dashboard. I just checked and the names I mentioned are not available right now, but if you put the cursor over their name and it says "so-and-so is solving some problem" you can click on the problem (in blue) and it will send you to the same page they are looking at. Then you can catch them while they're working and send them the link to your problem.

OpenStudy (anonymous):

alright, I don't want to intrude on someone else getting help either though lol But I will look out for them.

OpenStudy (turingtest):

If you post your link just say "can you look at this when you get the chance?" and they'll check it out after they're don with what they're doing.

OpenStudy (turingtest):

I just sent Zarkon this problem, hopefully he'll drop in soon.

OpenStudy (anonymous):

oh really? thanks!

OpenStudy (turingtest):

hmm... he looked at the problem and left. Hope he comes back...

OpenStudy (anonymous):

maybe he doesn't know either

OpenStudy (zarkon):

it's not correct as stated

OpenStudy (zarkon):

I think you need \[2^k-1\]

OpenStudy (anonymous):

oh yes you are right

OpenStudy (zarkon):

and \[\sum_{n=0}^{k-1}2^n=2^k-1\]

OpenStudy (anonymous):

yes I just noticed that I had put a 1 in there

OpenStudy (anonymous):

is there a way that I can break up that summation to show that it equals 2^k -1?

OpenStudy (zarkon):

this is a geometric sum..it follows \[\sum_{n=0}^{k-1}r^n=\frac{r^k-1}{r-1}\]

OpenStudy (anonymous):

I suppose there is no way to do it without the geometric sum formula. I haven't been taught that yet so I don't think I can use it.

OpenStudy (zarkon):

\[S=1+r+r^2+r^2+\cdots+r^{k-1}\] \[rS=r+r^2+r^2+\cdots+r^{k}\] subtract the two \[S-rS=1-r^k\] \[S(1-r)=1-r^k\] \[S=\frac{1-r^k}{1-r}=\frac{r^k-1}{r-1}\]

OpenStudy (zarkon):

gotta go...have fun

OpenStudy (anonymous):

thanks for your help

OpenStudy (turingtest):

Yeah he actually knows what he's talking about. One of the few...

OpenStudy (anonymous):

thanks for your help too!